# Dumbbell dimer dynamics in three-dimensional chiral fluids

### v1

@@ Abstract
We study the emergent orientational dynamics of a dumbbell dimer – two asymmetric monomers connected by a linking spring – in a three-dimensional chiral environment with odd viscosity. In classical systems with conserved parity symmetry, reciprocal oscillations of a dimer does not lead to rotational motion. Here, through an analytical calculation, we find that the presence of chirality in the system induces rotational dynamics as function of the expansion/contraction of the dimer. By incorporating thermal fluctuations, we further find that the rotational diffusivity is affected by the coupling between conformational fluctuations and rotational motion. Our results provide insights into problems where the parity symmetry is broken and can be used as a building block to study similar models at the collective level. These problems include multi-component molecular machines in odd-viscous fluids and systems with charged polymers where oddity is present through external magnetic fields.
@@ Introduction
Recently, the transport and dynamics of microparticles in chiral environments, which themselves break parity symmetries, has gained much attention and the macroscopic consequence of the self-spinning fluid components or particles in chiral motion has been studied [28, 29, 30]. Within the continuum level description, such fluid systems with broken parity exhibit intriguing chiral or “odd” responses, which are of current interest in the field [31, 32, 33]. From the symmetry argument, chiral systems naturally acquires an additional dissipationless transport coefficient called odd viscosity [34]. The emergence of the new quantitative measure has motivated and stimulated work proposing a new protocol to quantify active rheological properties experimentally [35]. The continuum description characterized by the odd viscosity has revealed various striking effects, such as transverse responses in the viscous, low-Reynolds-number regime [36, 37, 38, 39, 40, 41], as well as peculiar flocking and pattern formations observed at intermediate and high Reynolds numbers [42, 43]. Nevertheless, regardless of recent advances in theoretical studies, the hydrodynamic interactions and resulting multi-particle dynamics in chiral fluids have yet to be fully explored [44]. This is largely because exact analytical solutions of the fundamental hydrodynamic problem involving the response due to a point force and finite-sized particles remained scarce until just recently, especially in three-dimensional chiral fluids [45, 46], compared to two-dimensional counterparts [37, 47, 48, 49]. In two-dimension, the chiral systems are compatible with fluid isotropy [50] and can be relatively easily treated within the analytical framework using the complex notation [50, 36], whereas the emergent anisotropy leads to the nontrivial coupling between the spatial degrees of freedom in three-dimensional chiral systems.
@@
In this work, we study the dynamics of a dumbbell dimer immersed in unbounded three-dimensional chiral fluids. The dimer is either active by changing its conformation cyclically or passive, undergoing thermal fluctuations. Employing the continuum description of chiral systems characterized by odd viscosity, we derive the spatial equations of the dimer motion and show that the mobility of velocity-force relations obtains a non-symmetric structure in the exchange of spatial index. This suggests that odd viscosity naturally gives rise to an odd mobility in the spatial dynamics of particles and creates an intriguing rotational behavior, prohibited in conventional fluids. Using the mobility relations in fluids with odd viscosity, we study the two cases: the inter-dimer hydrodynamic interaction up to the leading order in odd viscosity and the higher order effect. Moreover, by extending the odd-viscous system in the presence of the thermal noise, we investigate the interplay between the fluctuation and oddity and how it affects the rotational dynamics of the dimer.
Model of a passive and active dimmer in viscous fluids
@@
We consider an axisymmetric dimer submerged in a three-dimensional incompressible chiral fluid with parity violation. This section recapitulates the concept of odd viscosity, a general measure that characterizes chiral fluid environments [31, 44]. At low Reynolds number, where inertial is negligible, the fluid property is governed by the force balance equation
@@
together with the incompressibility condition $\partial_{i}v_{i}=0$. Here $\bm{\sigma}$ represents the fluid the stress tensor, $\partial_{i}$ is the 3D differential operator with $i,j,k=\{x,y,z\}$, and the summation over repeated indices is assumed. The important relation here is the constitutive relation, which relates the stress field to the gradient of the velocity field:
@@
where $\eta_{ijk\ell}$ is the fourth-rank viscosity tensor and includes any fluid property. In order to introduce a minimum continuum description of chiral fluids, we assume the classical assumptions, such as the momentum conservation and the absence of the response to the vorticity, which are given by the equalities $\sigma_{ij}=\sigma_{ji}$ and $\eta_{ijk\ell}=\eta_{ij\ell k}$, respectively. To take into account the chirality in fluids, one should impose the left- and right-handedness on the viscosity tensor by introducing the parity violation around some axis, say, $\mathbf{e}$, while rotational invariance about the direction can still be assumed [34, 44]. Throughout this paper, we choose the axis of odd viscosity $\mathbf{e}$, without loss of generality, along the $z$-direction in the lab frame, i.e., $\mathbf{e}=\hat{\mathbf{z}}$. In this case, the viscosity tensor accounting for parity violation is simply expressed with two transport coefficients, shear and odd viscosities, $\eta^{\rm e}$ and $\eta^{\rm o}$ [34, 51]
@@
In an unbounded fluid described by the momentum conservation (1) and the viscosity tensor (Model of a passive and active dimmer in viscous fluids), we consider a dumbbell dimer composed of two spherical subunits with radius $a_{1}$ and $a_{2}$, connected via a shaft that can vary its length. The position of each sphere is denoted by the three-dimensional vector $\textbf{r}_{\alpha}$ ($\alpha=1,2$). Denoting the velocity of each sphere by $\dot{\textbf{r}}_{\alpha}=d\textbf{r}_{\alpha}/dt$ and the force acting on each subunit by $\textbf{f}_{\alpha}$, the overdamped equations of motion can be written as
@@
Here the mobility matrix $\textbf{M}_{\alpha\beta}$ accounts for the hydrodynamic contributions on the subunits exerted by the surrounding fluid. The diagonal mobility tensor $\textbf{M}_{\alpha\alpha}$ are associated with the hydrodynamic force acting on a each sphere, while the off-diagonal mobility tensor $\textbf{M}_{\alpha\beta}$ are related to the hydrodynamic interactions between the subunits. The off-diagonal components or Green’s function of the linear hydrodynamic equations further satisfy the relation $\partial_{j}M_{\alpha\beta,ij}=0$ from the incompressibility condition of the velocity field. In a conventional fluid with shear viscosity alone, theses two contributions are, respectively, attributed to Stokes’ law for a rigid sphere and the Oseen tensor, which is the solution of the forced Stokes equation [52]. In a fluid with odd viscosity, however, the chirality in the fluid gives rise to the anti-symmetric components in the mobility matrix, leading to asymmetric relations under the exchange of the spatial index, i.e., $(M_{\alpha\beta})_{ij}\neq(M_{\alpha\beta})_{ji}$ [47, 53, 54, 39, 41, 45, 38]. This condition holds for both the diagonal and off-diagonal mobility elements. Note that symmetric relations still hold under the particle notation, $\textbf{M}_{\alpha\beta}=\textbf{M}_{\beta\alpha}$.
This setup is generic and includes both the cases for passive and active dimers depending on the nature of the forces. The passive forces are typically gradient of conservative two-pair potentials minimally model through a harmonic potential. On the other hand, active forces could arise from non-equilibrium mechanochemical mechanisms that transduce chemical energy (typically through ATP hydrolysis) into mechanical active conformational oscillations [55, 56] which was recently reported in experiments of active nanomotors [57].
@@
For convenience, we introduce the relative frame of reference where $\textbf{L}=\textbf{r}_{2}-\textbf{r}_{1}$ and $\textbf{R}=(\textbf{r}_{1}+\textbf{r}_{2})/2$. Additionally, we write the relative vector as $\textbf{L}=L\textbf{n}$ where $L$ is the length of the dimer and n its orientation, which in spherical coordinates is expressed as $\textbf{n}=\sin\theta\cos\phi\hat{\textbf{x}}+\sin\theta\sin\phi\hat{\textbf{y}}+\cos\theta\hat{\textbf{z}}$ where $\theta$ is the polar angle measured from the $z$-axis (in the laboratory frame) and $\phi$ is the azimuthal angle around $\hat{\mathbf{z}}$. Since the considered active dimer autonomously changes its conformation without external forces, one should impose the force-free and torque-free conditions on it. This leads to the relations $_{2}=-\textbf{f}_{1}\equiv\textbf{f}$ f and $\textbf{n}\times\textbf{f}=\textbf{0}$. Starting from the velocity-force equations of each sphere to the leading order of odd viscosity, we obtain the equations of motion in the $(L_{i},R_{i})$ frame, which in the index notation are given by to the first order in the odd-to-even viscosity ratio $\lambda=\eta^{\rm o}/\eta^{\rm e}$
$\dot{L}_{i}=$ $\frac{f_{j}}{6\pi\eta^{\mathrm{e}}}\left(\frac{1}{a_{1}}+\frac{1}{a_{2}}\right)\left(\delta_{ij}-\frac{\lambda}{4}\epsilon_{ijz}\right)$
$-\frac{f_{j}}{4\pi\eta^{\mathrm{e}}L}\left(\delta_{ij}+n_{i}n_{j}-\frac{\lambda}{2}\epsilon_{ijz}\right),$ (5)
$\dot{R}_{i}=$ $\frac{f_{j}}{12\pi\eta^{\mathrm{e}}}\left(\frac{1}{a_{2}}-\frac{1}{a_{1}}\right)\left(\delta_{ij}-\frac{\lambda}{4}\epsilon_{ijz}\right).$ (6)
The first term in Eq. (5) results from $\eta^{\rm o}$-dependent Stokes’ law [39, 58], while the second represents the hydrodynamic interactions [54] and has been simplified by the torque-free condition. Equation (6) shows that the presence of odd viscosity can affect the center of mass motion when $a_{1}\neq a_{2}$. However, as stated by Purcell’s scallop theorem, dimer’s center of mass coordinate will be subjected only to a reciprocal motion over one cycle of the deformation [59]. We therefore focus on Eq. (5), in which the odd-mobility structure leads to an emergent rotational dynamics. Our goal is to obtain the dynamical equations for the orientational dynamics of the vector n for an arbitrary prescribed actuation of the dimer size $L(t)$.
@@
$\dot{\mathbf{L}}=\dot{L}\mathbf{n}+L\dot{\textbf{n}}.$ (7)
Through the evolution of the unit vector in terms of $(\theta,\phi)$, $\dot{\mathbf{n}}=\dot{\theta}\hat{\bm{\theta}}+\dot{\phi}\sin\theta\hat{\bm{\phi}}$ with $\hat{\bm{\theta}}=\cos\theta\cos\phi\hat{\textbf{x}}+\cos\theta\sin\phi\hat{\textbf{y}}-\sin\theta\hat{\textbf{z}}$ and $\hat{\bm{\phi}}=-\sin\phi\hat{\textbf{x}}+\cos\phi\hat{\textbf{y}}$, the dynamics of the two angles become
$\dot{\theta}$ $=L^{-1}(\dot{L}_{x}\cos\theta\cos\phi+\dot{L}_{y}\cos\theta\sin\phi-\dot{L}_{z}\sin\theta),$ (8)
$\dot{\phi}$ $=\frac{L_{x}\dot{L}_{y}-L_{y}\dot{L}_{x}}{(L\sin\theta)^{2}}.$ (9)
Applying the redefined torque-free conditions
$f_{x}$ $=\tan\theta\cos\phi f_{z},$ (10)
$f_{y}$ $=\tan\theta\sin\phi f_{z},$ (11)
to Eq. (5) and using Eqs. (8) and (9), we find that the polar angle evolution vanishes when $\lambda\ll 1$. However, the azimuthal angle evolution acquires a nonzero contribution, which is obtained as
$\dot{\phi}$ $=\frac{\lambda}{4}\frac{\dot{L}}{L}.$ (12)
Note that the obtained dynamics is independent of the azimuthal angle $\phi$, which is required by the system’s cylindrical symmetry around $\mathbf{e}$ [34]. To this order of the odd viscosity, the hydrodynamic interactions, which may be of linear order in $a_{\alpha}/L$, do not affect the orientational behavior and the angular dynamics is therefore independent of the domain size. It is obvious from the Eq. (12) that there is an emergence of the rotational dynamics due to the asymmetric mobility tensor, which couples the linear change in the interdomain distance $L(t)$ with motion transverse to it. In a classical fluid with shear viscosity alone or equivalently when $\lambda\to 0$, $\dot{\phi}$ vanishes, as it should. We note that for the special initial orientations, such that the dimer aligns with the axis of odd viscosity $(\theta=0,\pi)$, $\dot{\phi}$ vanishes. This is because Eqs. (10)-(11) vanish in this case and the anti-symmetric components driving the azimuthal rotation in Eq. (5) vanish as well. When the dimer orients in the plane perpendicular to the axis of odd viscosity ($\theta=\pi/2$), Eq. (12) still holds.
In the following, we provide an analytical solution of the dimer orientational dynamics without specifying the form of $L(t)$ The angular evolution (12) can be rewritten as
@@
For a non-vanishing odd viscosity, the dimer undergoes oscillatory rotation about $\hat{\textbf{z}}$ as a function of $L(t)$. When the dumbbell expands and the domains move apart, the dimer axis undergoes an transient anti-clockwise $(\lambda>0)$ or a clockwise precessional rotation $(\lambda<0)$, depending the sign of the odd viscosity. The results in this section suggest that at small values of odd viscosity, the angle azimuthal angle undergoes reciprocal oscillations as a function of $L(t)$, while the polar angle remains constant.

### v2

@@ Abstract
We study the emergent orientational dynamics of a dumbbell dimer – two asymmetric monomers connected by a linking spring – in a three-dimensional chiral environment with odd viscosity. In classical systems with conserved parity symmetry, reciprocal oscillations of a dimer do not lead to rotational motion. Here, through an analytical calculation, we find that the presence of chirality in the system induces rotational dynamics as a function of the expansion/contraction of the dimer. By incorporating thermal fluctuations, we further find that the rotational diffusivity is affected by the coupling between conformational fluctuations and rotational motion. Our results provide insights into problems where the parity symmetry is broken and can be used as a building block to study similar models at the collective level. These problems include multi-component molecular machines in odd-viscous fluids and systems with charged polymers where oddity is present through external magnetic fields.
@@ Introduction
Recently, the transport and dynamics of microparticles in chiral environments, which themselves break parity symmetries, has gained much attention and the macroscopic consequence of the self-spinning fluid components or particles in chiral motion has been studied [28, 29, 30]. Within the continuum level description, such fluid systems with broken parity exhibit intriguing chiral or “odd” responses, which are of current interest in the field [31, 32, 33]. From the symmetry argument, chiral systems naturally acquires an additional dissipationless transport coefficient called odd viscosity [34]. The emergence of the new quantitative measure has motivated and stimulated work proposing a new protocol to quantify active rheological properties experimentally [35]. The continuum description characterized by the odd viscosity has revealed various striking effects, such as transverse responses in the viscous, low-Reynolds-number regime [36, 37, 38, 39, 40, 41, 42], as well as peculiar flocking and pattern formations observed at intermediate and high Reynolds numbers [43, 44]. Nevertheless, regardless of recent advances in theoretical studies, the hydrodynamic interactions and resulting multi-particle dynamics in chiral fluids have yet to be fully explored [45]. This is largely because exact analytical solutions of the fundamental hydrodynamic problem involving the response due to a point force and finite-sized particles remained scarce until just recently, especially in three-dimensional chiral fluids [46, 47], compared to two-dimensional counterparts [37, 48, 49, 50]. In two-dimension, the chiral systems are compatible with fluid isotropy [51] and can be relatively easily treated within the analytical framework using the complex notation [51, 36], whereas the emergent anisotropy leads to the nontrivial coupling between the spatial degrees of freedom in three-dimensional chiral systems.
@@
In this work, we study the dynamics of a dumbbell dimer immersed in unbounded three-dimensional chiral fluids (Fig. 1). The dimer is either active by changing its conformation cyclically or passive, undergoing thermal fluctuations. Employing the continuum description of chiral systems characterized by odd viscosity, we derive the spatial equations of the dimer motion and show that the mobility of velocity-force relations obtains a non-symmetric structure in the exchange of spatial index. This suggests that odd viscosity naturally gives rise to an odd mobility in the spatial dynamics of particles and creates an intriguing rotational behavior [Fig. 1(b,c)], prohibited in conventional fluids. Using the mobility relations in fluids with odd viscosity, we study the two cases: the interdomain hydrodynamic interaction up to the leading order in odd viscosity and the higher order effect. Moreover, by extending the odd-viscous system in the presence of the thermal noise, we investigate the interplay between the fluctuation and oddity and how it affects the rotational dynamics of the dimer.
Figure 1: Schematic of a dumbbell dimer immersed in a three-dimensional chiral fluid with an even (shear) viscosity and an odd viscosity, $\eta^{\rm e}$ and $\eta^{\mathrm{o}}$. The odd viscosity is assumed to be positive. The axis of the odd viscosity is assumed to be in the $\hat{\textbf{z}}-$direction, while the even viscosity is isotropic and thus does not have any preferred direction in the fluid. (a) The geometry of a dimer whose spherical domains have radii $a_{1}$ and $a_{2}$. The dimer length as well as the dimer orientation vary in time, which are characterized by the scalar function $L(t)$ and the unit vector $\textbf{n}(t)$, respectively. The dimer length undergoes expansions or contractions due to thermal noise or prescribed motion of the connecting shaft, which leads to the three dimensional rotational motion of the dimer. (b) To linear order in the odd viscosity, the expansion of the dimer arm creates an anti-clockwise rotation about the $\hat{\textbf{z}}-$axis by increasing the azimuthal angle $\phi(t)$ measured from the $x$-axis [see also Eq. (17) and Fig. 2(b)]. (c) To quadratic order in the odd viscosity, in addition to the transient precession observed in panel (b), the contraction of the dimer causes the reorientation towards the $(x,y)$ plane by increasing the polar angle $\theta(t)$ measured from the $z$-axis [see also Eq. (22) and Fig. 2(a)].
Model of a passive and active dimer in viscous fluids
@@
We consider an axisymmetric dimer submerged in a three-dimensional incompressible chiral fluid with parity violation (Fig. 1). This section recapitulates the concept of odd viscosity, a general measure that characterizes chiral fluid environments [31, 45]. At low Reynolds number, where inertial is negligible, the fluid property is governed by the force balance equation
@@
together with the incompressibility condition $\partial_{i}v_{i}=0$. Here $\bm{\sigma}$ represents the fluid stress tensor, $\partial_{i}$ is the three-dimensional differential operator with $i,j,k=\{x,y,z\}$, and the summation over repeated indices is assumed. The important relation here is the constitutive relation, which relates the stress field to the gradient of the velocity field:
@@
where $\eta_{ijk\ell}$ is the fourth-rank viscosity tensor and includes any fluid property. In order to introduce a minimum continuum description of chiral fluids, we assume the classical assumptions, such as the momentum conservation and the absence of the response to the vorticity, which are given by the equalities $\sigma_{ij}=\sigma_{ji}$ and $\eta_{ijk\ell}=\eta_{ij\ell k}$, respectively. To take into account the chirality in fluids, one should impose the left- and right-handedness on the viscosity tensor by introducing the parity violation around some axis, while rotational invariance about the direction can still be assumed [34, 45]. Throughout this paper, we choose the axis of odd viscosity along the $z$-direction in the lab frame, without loss of generality. In this case, the viscosity tensor accounting for parity violation is simply expressed with two transport coefficients, shear and odd viscosities, $\eta^{\rm e}$ and $\eta^{\rm o}$ [34, 52]
@@
In an unbounded fluid described by the momentum conservation (1) and the viscosity tensor (Model of a passive and active dimer in viscous fluids), we consider a dumbbell dimer composed of two spherical subunits with radius $a_{1}$ and $a_{2}$, connected via a shaft that can vary its length. The position of each sphere is denoted by the three-dimensional vector $\textbf{r}_{\alpha}$ ($\alpha=1,2$). Denoting the velocity of each sphere by $\dot{\textbf{r}}_{\alpha}=d\textbf{r}_{\alpha}/dt$ and the force acting on each subunit by $\textbf{f}_{\alpha}$, the overdamped equations of motion can be written as
@@
Here the mobility matrix $\textbf{M}_{\alpha\beta}$ accounts for the hydrodynamic contributions on the subunits exerted by the surrounding fluid. The diagonal mobility tensor $\textbf{M}_{\alpha\alpha}$ is associated with the hydrodynamic force acting on a each sphere, while the off-diagonal mobility tensor $\textbf{M}_{\alpha\beta}$ is related to the hydrodynamic interactions between the subunits. The off-diagonal components or Green’s function of the linear hydrodynamic equations further satisfy the relation $\partial_{j}M_{\alpha\beta,ij}=0$ from the incompressibility condition of the velocity field. In a conventional fluid with shear viscosity alone, theses two contributions are, respectively, attributed to Stokes’ law for a rigid sphere and the Oseen tensor, which is the solution of the forced Stokes equation [53]. In a fluid with odd viscosity, however, the chirality in the fluid gives rise to the anti-symmetric components in the mobility matrix, leading to asymmetric relations under the exchange of the spatial index, i.e., $(M_{\alpha\beta})_{ij}\neq(M_{\alpha\beta})_{ji}$ [48, 54, 55, 39, 41, 46, 38]. This condition holds for both the diagonal and off-diagonal mobility elements. Note that symmetric relations still hold under the particle notation, $\textbf{M}_{\alpha\beta}=\textbf{M}_{\beta\alpha}$.
When the distances between the spheres are sufficiently larger than their sizes $(a_{\alpha}\ll|\mathbf{r}_{\alpha}-\mathbf{r}_{\beta}|\equiv r_{\alpha\beta})$, $\mathbf{M}_{\alpha\beta}$ can be approximated as [46]
[…]
$\left.+\left(\frac{m_{+}\lambda^{2}}{2}+4\right)\hat{\mathbf{z}}\hat{\mathbf{z}}+f\lambda\bm{\epsilon}\cdot\hat{\mathbf{z}}\right$], (5)
$\mathbf{M}_{\alpha\beta}=$ $\frac{1}{4\pi\eta^{\rm e}(r_{\alpha\beta}+r_{\lambda})}$
$\times\Bigg[\mathbf{I}+\mathbf{nn}-\frac{\lambda r_{\alpha\beta}}{2r_{\lambda}}\bm{\epsilon}\cdot\left(\hat{\mathbf{z}}-(\mathbf{n}\cdot\hat{\mathbf{z}})\mathbf{n}\right)$
$-\left(1-\frac{r_{\alpha\beta}}{r_{\lambda}}\right)\left(\mathbf{nn}+\frac{(\mathbf{n}\times\hat{\mathbf{z}})(\mathbf{n}\times\hat{\mathbf{z}})}{1-(\textbf{n}\cdot\hat{\textbf{z}})^{2}}\right)\Bigg],$ (6)
where $r_{\lambda}=r_{\alpha\beta}\sqrt{4+\lambda^{2}[1-(\textbf{n}\cdot\hat{\textbf{z}})^{2}]}/2$ and $\textbf{n}=(\textbf{r}_{\alpha}-\textbf{r}_{\beta})/r_{\alpha\beta}$ denotes a unit vector along the axis of a dumbbell dimer [Fig. 1(a)]. In the above, the scaling functions of $\lambda$ are defined as $m_{+}=f+g$ and $m_{-}=f-g$ with $f=-12\left(2\psi+\lambda\right)\lambda^{-3}$, $g=4(1+f)\lambda^{-2}$, and $\psi=\arcsin\bigl(-\lambda(4+\lambda^{2})^{-1/2}\bigr)$ [46]. Expression (5) corresponds to the odd Stokes mobility, while Eq. (6) describes the hydrodynamic interaction governed by Green’s function, which is the solution of a point force acting on an unbounded fluid with $\eta^{\rm e}$ and $\eta^{\rm o}$. These expressions are valid for any odd-to-even viscosity ratio, $\lambda=\eta^{\rm o}/\eta^{\rm e}$. Note that the hydrodynamic effects on the subunits due to the presence of the connecting shaft are assumed to be negligible [56], so that the interaction can be described by the Green’s function of a point force, as in Eq. (6).
This setup is generic and includes both the cases for passive and active dimers depending on the nature of the forces. Passive forces are typically the gradient of a conservative two-pair potential minimally modeled by a harmonic potential. On the other hand, active forces could arise from non-equilibrium mechanochemical mechanisms that transduce chemical energy (typically through ATP hydrolysis) into mechanical active conformational oscillations [57, 58] that was recently reported in experiments of active nanomotors [59].
@@
To determined the dynamics of an active dumbbell dimer, it is convenient to introduce the relative frame of reference where $\textbf{L}=\textbf{r}_{2}-\textbf{r}_{1}$ and $\textbf{R}=(\textbf{r}_{1}+\textbf{r}_{2})/2$. Additionally, we write the relative vector as $\textbf{L}=L\textbf{n}$ where $L$ is the length of the dimer and n its orientation. In spherical coordinates, they are expressed as $\textbf{n}=\sin\theta\cos\phi\hat{\textbf{x}}+\sin\theta\sin\phi\hat{\textbf{y}}+\cos\theta\hat{\textbf{z}}$ where $\theta$ is the polar angle measured from the $z$-axis (in the laboratory frame) and $\phi$ is the azimuthal angle around $\hat{\mathbf{z}}$ (see Fig. 1). Unless otherwise stated, we consider situations in which dimers deform their shapes in a time-reversal manner, i.e., undergoing reciprocal motion. In general, however, the dimer model is not limited to this type of deformation and can also exhibit any deformation specified by the dimer length, $L(t)$. Since the considered active dimer autonomously changes its conformation without external forces, one should impose the force-free and torque-free conditions on it. This leads to the relations $\textbf{F}_{2}=-\textbf{F}_{1}\equiv\textbf{F}$ and $\textbf{n}\times\textbf{F}=\textbf{0}$.
By using the relations $\dot{\mathbf{L}}=(\mathbf{M}_{11}+\mathbf{M}_{22}-2\mathbf{M}_{12})\cdot\mathbf{F}$ and $\dot{\mathbf{R}}=(\mathbf{M}_{22}-\mathbf{M}_{11})\cdot\mathbf{F}/2$ and applying the torque-free condition to them, the dynamics of the dimer conformation simplifies to
$\dot{\mathbf{L}}$ $=\mathbf{M}_{+}\cdot\mathbf{F}$
$-\frac{1}{2\pi\eta^{\rm e}(r_{12}+r_{\lambda})}\left[\left(1+\frac{r_{12}}{r_{\lambda}}\right)\mathbf{I}-\frac{\lambda r_{12}}{2r_{\lambda}}\bm{\epsilon}\cdot\hat{\mathbf{z}}\right]\cdot\mathbf{F},$ (7)
$\dot{\mathbf{R}}$ $=-\frac{1}{2}\mathbf{M}_{-}\cdot\mathbf{F},$ (8)
where we have defined the self-mobility difference as
$\mathbf{M}_{\pm}=\frac{1}{24\pi\eta^{\rm e}}\left(\frac{1}{a_{1}}\pm\frac{1}{a_{2}}\right)$
$\times\left[\biggl(\frac{m_{-}\lambda^{2}}{4}+4\biggr)\mathbf{I}+\frac{\lambda^{2}}{2}\biggl(m_{+}-\frac{m_{-}}{2}\biggr)\hat{\mathbf{z}}\hat{\mathbf{z}}+f\lambda\bm{\epsilon}\cdot\hat{\mathbf{z}}\right$]. (9)
In the presence of odd viscosity $(\eta^{\rm o}\neq 0)$, the self-mobility causes velocity not only along the dimer director $\mathbf{n}$, but also in the $z$- and azimuthal directions. Notice that the transverse force along $\hat{\mathbf{z}}$ is absent in the hydrodynamic interaction described by the Green’s function.
@@
Starting from the velocity-force equations (Orientational dynamics of active dimer) and (8) and expanding them to the first order in $\lambda=\eta^{\rm o}/\eta^{\rm e}$, we obtain the equations of motion, which are given by
[…]
$-\frac{1}{{\color[rgb]{0,0,0}2}\pi\eta^{\mathrm{e}}L}\left({\color[rgb]{0,0,0}\textbf{I}}-\frac{\lambda}{{\color[rgb]{0,0,0}4}}\bm{\epsilon}\cdot\hat{\textbf{z}}\right)\Bigg]\cdot\textbf{F},$ (10)
$\dot{\textbf{R}}=$ $\frac{1}{12\pi\eta^{\mathrm{e}}}\left(\frac{1}{a_{2}}-\frac{1}{a_{1}}\right)\left(\textbf{I}-\frac{\lambda}{4}\bm{\epsilon}\cdot\hat{\textbf{z}}\right)\cdot\textbf{F}.$ (11)
The first term in Eq. (10) results from $\eta^{\rm o}$-dependent Stokes’ law [39, 60], while the second represents the hydrodynamic interactions [55] and has been simplified by the torque-free condition. Equation (11) shows that the presence of odd viscosity can affect the center of mass motion when $a_{1}\neq a_{2}$. However, dimer’s center of mass coordinate will be subject only to an oscillatory motion without net displacement over one cycle of the reciprocal deformation. During such conformational changes, the dimers deform its shapes in a time-reversal manner, and therefore self-propelled motion does not occur, as in the case of classical fluids. This is known as Purcell’s scallop theorem [61]. We therefore focus on Eq. (10), in which the odd-mobility structure leads to an emergent rotational dynamics. Our goal is to obtain the dynamical equations for the orientational dynamics of the vector $\textbf{n}(t)$ for an arbitrary prescribed actuation of the dimer size $L(t)$.
@@
$\dot{\mathbf{L}}=\dot{L}\mathbf{n}+L\dot{\textbf{n}},$ (12)
where the evolution of the unit vector is expressed in terms of the polar and the azimuthal angles, $\theta,\phi$, as $\dot{\mathbf{n}}=\dot{\theta}\hat{\bm{\theta}}+\dot{\phi}\sin\theta\hat{\bm{\phi}}$ with $\hat{\bm{\theta}}=\cos\theta\cos\phi\hat{\textbf{x}}+\cos\theta\sin\phi\hat{\textbf{y}}-\sin\theta\hat{\textbf{z}}$ and $\hat{\bm{\phi}}=-\sin\phi\hat{\textbf{x}}+\cos\phi\hat{\textbf{y}}$. In the coordinate system with the orthogonal bases $\{\mathbf{n},\hat{\bm{\theta}},\hat{\bm{\phi}}\}$, one finds $\dot{\textbf{L}}=(\dot{L},L\dot{\theta},L\dot{\phi}\sin\theta)$, and the torque-free condition is simply given by $\hat{\bm{\theta}}\cdot\textbf{F}=\hat{\bm{\phi}}\cdot\textbf{F}=0$. We then find $\textbf{F}=(F_{n},0,0)$ with $F_{n}=\textbf{n}\cdot\textbf{F}$. Expressing the velocity-force relation (10) in the spherical coordinates, we find that the polar angle evolution vanishes ($\dot{\theta}=0$), while the dynamics of the length and the azimuthal angle are nonzero:
$\dot{L}$ $=\left[\frac{1}{6\pi\eta^{\mathrm{e}}}\left(\frac{1}{a_{1}}+\frac{1}{a_{2}}\right)-\frac{1}{2\pi\eta^{\mathrm{e}}L}\right]F_{n},$ (13)
$\dot{\phi}$ $=\frac{\lambda}{4L}\left[\frac{1}{6\pi\eta^{\mathrm{e}}}\left(\frac{1}{a_{1}}+\frac{1}{a_{2}}\right)-\frac{1}{2\pi\eta^{\mathrm{e}}L}\right]F_{n}.$ (14)
Comparing Eqs. (13) and (14) simply gives
$\dot{\phi}$ $=\frac{\lambda}{4}\frac{\dot{L}}{L},$ (15)
for $0<\theta<\pi$. Note here that for the special initial orientations when the dimer aligns with the axis of odd viscosity $(\theta=0,\pi)$, the dimer does not exhibit any rotational motion $(\dot{\phi}=0)$. This is because the anti-symmetric components driving the azimuthal rotation vanish when $\theta=0,\pi$ [see Eq. (10)]. Moreover, the obtained dynamics is independent of the azimuthal angle $\phi$, which is required by the system’s cylindrical symmetry around $\hat{\mathbf{z}}$ [34]. Moreover, to this order, the dynamics of the azimuthal angle are proportional to the dynamics of the conformational speed of the dimer length. It is obvious from the Eq. (15) that there is an emergence of the rotational dynamics due to the asymmetric mobility tensor, which couples the linear change in the interdomain distance $L(t)$ with motion transverse to it. In a classical fluid with shear viscosity alone or equivalently when $\lambda\to 0$, $\dot{\phi}$ vanishes, as it should. When the dimer orients in the plane perpendicular to the axis of odd viscosity ($\theta=\pi/2$), Eq. (15) still holds.
In the following, we provide an analytical solution of the dimer orientational dynamics without specifying the form of $L(t)$. The angular evolution (15) can be rewritten as
@@
For a non-vanishing odd viscosity, the dimer undergoes oscillatory rotation about $\hat{\textbf{z}}$ as a function of $L(t)$. When the dumbbell expands and the domains move apart, the dimer axis undergoes a transient anti-clockwise $(\lambda>0)$ [Fig. 1(b)] or a clockwise precessional rotation $(\lambda<0)$, depending the sign of the odd viscosity. The results in this section suggest that at small values of odd viscosity, the azimuthal angle exhibits the oscillatory behavior as a function of $L(t)$, while the polar angle remains constant.